# Attention-Aided MMSE for OFDM Channel Estimation: Learning Linear Filters with Attention

### v3

@@ III-B DNN-based Approach
To address the limitations of SP-based approaches, DNN-based approaches have been actively explored as a data-driven alternative. The key idea behind existing DNN-based methods is to leverage DNNs as a denoiser and an interpolator for channel estimation. Specifically, by modeling the DNN input as the noisy received pilot signal (i.e., ${\bf{Y}}_{p}$) and the output as the corresponding channel estimate (i.e., $\hat{\bf{H}}$), we train the DNN to learn a nonlinear mapping between ${\bf{Y}}_{p}$ and $\hat{\bf{H}}$. Existing research on DNN-based approaches can be categorized according to the specific neural network architectures employed for channel estimation.
ChannelNet: ChannelNet [soltani:commlett:19] represents one of the initial frameworks that exploits a DNN for OFDM channel estimation. In this approach, the channel response is treated as an image. An initial estimate of the channel is obtained by applying bilinear interpolation to sparsely located pilot symbols, resulting in a low-resolution and noise-contaminated channel map. Then, ChannelNet uses a super-resolution convolutional neural network (i.e., SRCNN) to recover high-resolution spatial features from the interpolated input. After this, a denoising convolutional neural network (i.e., DnCNN) is applied, incorporating residual learning and batch normalization to further refine the estimate.
Channelformer: Channelformer [luan2023channelformer] exploits the Attention Transformer architecture for OFDM channel estimation. Unlike ChannelNet, Channelformer flattens the LS channel estimate for all pilot subcarriers into a single vector and feeds its real and imaginary parts as two separate input channels. Based on this, the encoder captures long-range dependencies among sparse pilot embeddings using the Attention Transformer architecture. Instead of the standard Feed-Forward Network, it employs convolutional layers to introduce local inductive bias, while retaining global modeling capabilities through multi-head self-Attention. The decoder refines these context-aware features using a residual convolutional network, integrating both global Attention and local structural information.
For better understanding, we explain the principle of Channelformer as follows. Fundamentally, Channelformer follows the Multi-Head Self-Attention (MHA) mechanism [vaswani2017attention]. To be specific, denoting $p_{i}$ as the complex-valued channel observation at pilot index $i$ with $1\leq i\leq L$, each pilot observation is first embedded into a two-dimensional real-valued row vector that retains its inherent complex structure: $\mathbf{e}_{i}=\left[\Re(p_{i}),\Im(p_{i})\right]\in\mathbb{R}^{2}$. Then, we define the full sequence of embedded pilot vectors as
@@
where the $h$ is equal to 2, representing the number of symbols containing pilots and the Attention operation within each head is the Scaled Dot-Product Attention defined as
@@ IV-A A-MMSE Estimation Mechanism
where ${\bf{W}}_{\text{A-MMSE}}$ represents the A-MMSE linear filter. The A-MMSE filter ${\bf{W}}_{\text{A-MMSE}}$ is learned through our Attention Transformer architecture as follow:
@@
It is important to highlight several key points regarding the proposed A-MMSE method. First, by leveraging the Attention Transformer to learn ${\bf{W}}_{\text{A-MMSE}}$, A-MMSE inherits the primary advantages of DNN-based approaches: powerful learning capability and no requirement for prior channel statistics. Second, once ${\bf{W}}_{\text{A-MMSE}}$ is learned, we only use linear computation as presented in (19), avoiding the computational complexity associated with nonlinear activation functions in the channel estimation process. In this sense, the proposed A-MMSE is a hybrid approach that effectively combines SP-based and DNN-based approaches. Further, in the A-MMSE framework, it is possible to adjust the number of effective parameters within ${\bf{W}}_{\text{A-MMSE}}$, so that the A-MMSE adapts its computational overheads depending on various environments. We further explore this in the next section.
@@ IV-B1 Insights on Two-Stage Attention Encoder
We now provide a detailed explanation of the A-MMSE learning mechanism $\mathcal{F}_{\text{A-MMSE}}(\cdot)$ (IV-A), whose overall architecture and training pipeline are depicted in Fig. 2. Structurally inspired by recent encoder-decoder architectures based on Attention Transformer, the proposed A-MMSE method incorporates theoretical insights derived from classical SP-based channel estimation approaches. Specifically, as shown in the MMSE filter structure (9), the optimal linear filter operates by first whitening the pilot observations and then interpolating the full OFDM channel using the underlying frequency-temporal domain channel correlation. This highlights the importance of explicitly incorporating the channel’s second-order statistics. The proposed A-MMSE learning mechanism effectively captures the correlation structure inherent to OFDM channels. Its detailed formulation is presented as follows.
Under the assumption of a wide-sense stationary uncorrelated scattering (WSSUS) environment and separable scattering function [Yoojin20152DLMMSE], the second-order statistics of OFDM channels exhibit a separable correlation function across frequency and temporal domains. Formally, the channel correlation matrix between two channel taps separated by frequency index difference $\Delta n$ and time index difference $\Delta m$ is factorized as follows:
@@
where $\mathbf{h}$ denotes $\operatorname{vec}(\mathbf{H})\in\mathbb{C}^{NM}$, $\mathbf{R}_{f}\in\mathbb{C}^{N\times N}$ and $\mathbf{R}_{t}\in\mathbb{C}^{M\times M}$ represent the frequency and temporal domain correlation matrices, respectively, as visualized in Fig. 3. A detailed proof of this Kronecker product structure is provided in Appendix A. We note that the approximation ($\approx$) in (22) arises because the WSSUS assumption may not perfectly hold in practical scenarios. The representation of the full covariance matrix ${\bf{R}}_{\text{full}}$ as the Kronecker product ($\otimes$) between the frequency-domain correlation ${\bf{R}}_{f}$ and the temporal-domain correlation ${\bf{R}}_{t}$ (22) implies that the frequency-temporal domain correlation structure is separable. This allows for a compact and computationally efficient representation of the channel covariance matrix. Building on this insight, we propose the following two-stage encoder design.
  • Frequency Encoder: In the Frequency Encoder stage, the pilot observations are processed along the subcarrier domain, extracting features on the frequency-domain correlation structure at pilot indices, corresponding to row of the matrix $\mathbf{R}_{f}\left[\mathcal{P},:\right]$.
  • Temporal Encoder: In the Temporal Encoder stage, the output of the Frequency Encoder is then subsequently processed along the OFDM symbol domain, capturing temporal-domain correlation structures at pilot indices, corresponding to rows of the matrix $\mathbf{R}_{t}\left[\mathcal{P},:\right]$.
This two-stage encoder enables the model to learn the joint frequency–temporal correlation structure; each stage uses a single encoder block, which is critical for the accurate MMSE estimation. Crucially, we note that this encoder design represents a key distinction of the proposed A-MMSE learning mechanism compared to conventional DNN-based methods. Specifically, existing approaches typically adopt standard learning architectures such as CNN [soltani:commlett:19] or Attention Transformer [luan2023channelformer], without explicitly tailoring them to the statistical properties inherent in OFDM channels. In contrast, our method explicitly incorporates the inherent temporal and frequency-domain correlation structures of OFDM channels into the learning process. This integration not only enhances learning efficiency but also improves the model’s capability to generalize under practical channel conditions. In this sense, the proposed A-MMSE approach is distinguished not only by its design, which enables the learning of a linear estimation filter rather than directly estimating the OFDM channel matrix (as illustrated in Fig.1), but also by its advanced learning mechanism that explicitly reflects the inherent correlation structure of OFDM channels.
@@ IV-B2 Multi-Head Self-Attention Mechanism
Subsequently, the embedding matrix $\mathbf{X}_{\text{Emb}}$ is fed into the MHA module $\left(\mathcal{F}_{\text{MHA}}(\cdot)\right)$ with the number of heads fixed at $L$
, where $L$ equals the total number of DM-RS pilot symbol (one head per pilot). Each complete pilot is split into real and imaginary parts, resulting in the $2L$ column dimension of $\mathbf{X}_{\mathrm{Emb}}$. For example, with Type-A DM-RS and additional Position =1=1 (i.e., two DM-RS OFDM symbols) using a comb-2 pattern over 6 reousrce blocks (RBs), that has 72 subcarriers, there are $72/2=36$ pilots per DM-RS symbol, hence $L=36\times 2=72$ pilots in total. In general,
@@
where $N_{\mathrm{RB}}$ is the number of resource blocks, $N_{\mathrm{SC/RB}}=12$ is the number of subcarriers per RB, and $N_{\mathrm{DM-RS\,sym}}$ denotes the number of OFDM symbols carrying DM-RS. The operation of the MHA module $\mathcal{F}_{\text{MHA}}(\cdot)$ is described in (13)–(16). We note that this allows each head to learn distinct frequency domain correlation patterns. Following the MHA module, Add & Layer Normalization is applied, succeeded by a FFN, and subsequently another Add & Layer Normalization, in sequential order. Details on Add & Norm and FFN are provided in [vaswani2017attention]. After these processes, the encoder extracts frequency-domain statistical information, which can be formally represented as
@@
where $\mathbf{Y}_{\text{Freq}}\in\mathbb{R}^{2L\times N}$ represents features encoded by the Frequency Encoder, $\mathcal{E}_{\text{Frequency}}$ represents the Frequency Encoder module, and $\theta_{\text{MHA, Freq}}$ and $\theta_{\text{Freq}}$ denote the learnable parameters of MHA module and the combined parameters of the two Add & Layer Normalization and FFN, respectively. Correspondingly, the overall operation $\mathbf{x}_{\text{in}}\mapsto\mathbf{Y}_{\text{Freq}}$ yields a real-valued matrix, where each row corresponds to a pilot position and columns aligns with subcarrier-specific embedding dimensions, explicitly capturing both local and global dependencies in the frequency domain.
Temporal Encoder: In the second stage, a linear projection layer transforms the features encoded in the previous stage into a high dimensional frequency-temporal domain, where the projection dimension is $NM$. The encoded feature matrix is given as $\mathbf{Y}_{\text{Freq}}=\left[\mathbf{y}_{\text{freq},1};\mathbf{y}_{\text{freq},2};\cdots;\mathbf{y}_{\text{freq},2L}\right]\in\mathbb{R}^{2L\times N}$. Here, each row vector $\mathbf{y}_{\text{Freq},i}$ is linearly projected as $\mathbf{x}_{\text{proj},i}=\mathbf{y}_{\text{Freq},i}\cdot\mathbf{P}_{\text{proj}}+\mathbf{b}_{\text{proj}}$, where $\mathbf{P}_{\text{proj}}\in\mathbb{R}^{N\times d_{p}}$ is a learnable projection matrix and $\mathbf{b}_{\text{proj}}\in\mathbb{R}^{d_{p}}$ is a learnable bias vector. Stacking all projected vectors $\mathbf{x}_{\text{proj},i}$, this yields the matrix given by
@@
Next, the projected matrix $\mathbf{X}_{\text{Proj}}$ is fed into another MHA module, with the number of heads fixed at $M$ (one head for each OFDM symbol). Similar to the above, this enables each to effectively learn distinct temporal correlation patterns. The subsequent process is identical to that of the Frequency Encoder.
@@
where $\mathbf{Y}_{\text{Temp}}\in\mathbb{R}^{2L\times NM}$ represents features encoded by the Temporal Encoder, $\mathcal{E}_{\text{Temporal}}$ represents the Temporal Encoder module, with $\theta_{\text{MHA, Temp}}$ and $\theta_{\text{Temp}}$ being the learnable parameters of MHA module and the combined parameters of the two Add & Layer Normalization and FFN, respectively. The overall operations $\mathbf{Y}_{\text{Freq}}\mapsto\mathbf{Y}_{\text{Temp}}$ yield a real-valued matrix, whose rows correspond to pilot positions and whose columns align with subcarrier-symbol specific embedding dimensions. This explicitly captures both local and global dependencies in the frequency-temporal domain of the OFDM channel. The frequency-temporal features encoded by the two-stage encoder serve as inputs to the decoder.
@@
Once $\mathbf{Y}_{\text{Temp}}$ is obtained, it is partitioned into the real and imaginary components as follows.
@@ IV-B3 Training Method
where $N_{\mathcal{B}}$ denotes the number of mini batch and $\mathcal{B}$ denotes the mini batch size. In addition, $\mathbf{H}_{i,j}$ and $\hat{\bf{H}}_{i,j}$ represent the corresponding true channel matrix and estimated channel matrix, respectively. For the loss function $\mathcal{L}(\cdot)$, MSE or Huber loss can be chosen. We discuss this further in the next subsection. The loss function (32) is minimized via backpropagation and the Adam optimizer [kingma2014adam].

### v4

@@ III-B DNN-based Approach
To address the limitations of SP-based approaches, DNN-based approaches have been actively explored as a data-driven alternative. The key idea behind existing DNN-based methods is to leverage DNNs as a denoiser and also an interpolator for channel estimation. Specifically, by modeling the DNN input as the noisy received pilot signal (i.e., ${\bf{Y}}_{p}$) and the output as the corresponding channel estimate (i.e., $\hat{\bf{H}}$), we train the DNN to learn a nonlinear mapping between ${\bf{Y}}_{p}$ and $\hat{\bf{H}}$. Existing research on DNN-based approaches can be categorized according to the specific neural network architectures employed for channel estimation.
ChannelNet: The ChannelNet [36] represents one of the initial frameworks that exploits a DNN for OFDM channel estimation. In this approach, the channel response is treated as an image. An initial estimate of the channel is obtained by applying bilinear interpolation to sparsely located pilot symbols, resulting in a low-resolution and noise-contaminated channel map. Then, ChannelNet uses a super-resolution convolutional neural network (i.e., SRCNN) to recover high-resolution spatial features from the interpolated input. After this, a denoising convolutional neural network (i.e., DnCNN) is applied, incorporating residual learning and batch normalization to further refine the estimate.
Channelformer: The Channelformer [25] exploits the Attention Transformer architecture for OFDM channel estimation. Unlike ChannelNet, Channelformer flattens the LS channel estimate for all pilot subcarriers into a single vector and feeds its real and imaginary parts as two separate input channels. Based on this, the encoder captures long-range dependencies among sparse pilot embeddings using the Attention Transformer architecture. Instead of the standard Feed-Forward Network, it employs convolutional layers to introduce local inductive bias, while retaining global modeling capabilities through multi-head self-Attention. The decoder refines these context-aware features using a residual convolutional network, integrating both global Attention and local structural information.
To clarify the mechanism, we explain the principle of Channelformer as follows. Fundamentally, Channelformer follows the Multi-Head Self-Attention (MHA) mechanism [37]. To be specific, denoting $p_{i}$ as the complex-valued channel observation at pilot index $i$ with $1\leq i\leq L$, each pilot observation is first embedded into a two-dimensional real-valued row vector that retains its inherent complex structure: $\mathbf{e}_{i}=\left[\Re(p_{i}),\Im(p_{i})\right]\in\mathbb{R}^{2}$. Then, we define the full sequence of embedded pilot vectors as
@@
where $h=2$, representing the number of symbols containing pilots and the Attention operation within each head is the Scaled Dot-Product Attention defined as
@@ IV-A A-MMSE Estimation Mechanism
where ${\bf{W}}_{\text{A-MMSE}}$ represents the A-MMSE linear filter. The A-MMSE filter ${\bf{W}}_{\text{A-MMSE}}$ is learned through our Attention Transformer architecture as follows:
@@
It is important to highlight several key points regarding the proposed A-MMSE method. First, by leveraging the Attention Transformer to learn ${\bf{W}}_{\text{A-MMSE}}$, A-MMSE inherits the primary advantages of DNN-based approaches: powerful learning capability and no requirement for prior channel statistics. Second, once ${\bf{W}}_{\text{A-MMSE}}$ is learned, we only use linear computation as presented in (18), avoiding the computational complexity associated with nonlinear activation functions in the channel estimation process. In this sense, the proposed A-MMSE is a hybrid approach that effectively combines SP-based and DNN-based approaches. Further, in the A-MMSE framework, it is possible to adjust the number of effective parameters within ${\bf{W}}_{\text{A-MMSE}}$, enabling the A-MMSE to adapt its computational overheads depending on various environments. We further explore this in the next section.
@@ IV-B1 Insights on Two-Stage Attention Encoder
We now provide a detailed explanation of the A-MMSE learning mechanism $\mathcal{F}_{\text{A-MMSE}}(\cdot)$ (IV-A), whose overall architecture and training pipeline are depicted in Fig. 2. Structurally inspired by recent encoder-decoder architectures based on Attention Transformer, the proposed A-MMSE method incorporates theoretical insights derived from classical SP-based channel estimation approaches. Specifically, as shown in the MMSE filter structure (9), the optimal linear filter operates by first whitening the pilot observations and then interpolating the full OFDM channel using the underlying frequency-temporal-domain channel correlation. This highlights the importance of explicitly incorporating the channel’s second-order statistics. The proposed A-MMSE learning mechanism effectively captures the correlation structure inherent to OFDM channels. Its detailed formulation is presented as follows.
Under the assumption of a wide-sense stationary uncorrelated scattering (WSSUS) environment and separable scattering function [5], the second-order statistics of OFDM channels exhibit a separable correlation function across frequency and temporal domains. Formally, the channel correlation matrix between two channel taps separated by frequency index difference $\Delta n$ and time index difference $\Delta m$ is factorized as follows:
@@
where $\mathbf{h}$ denotes $\operatorname{vec}(\mathbf{H})\in\mathbb{C}^{NM}$, $\mathbf{R}_{f}\in\mathbb{C}^{N\times N}$ and $\mathbf{R}_{t}\in\mathbb{C}^{M\times M}$ represent the frequency and temporal domain correlation matrices, respectively, as visualized in Fig. 3. We note that the approximation ($\approx$) in (21) arises because the WSSUS assumption may not perfectly hold in practical scenarios. The representation of the full covariance matrix ${\bf{R}}_{\text{full}}$ as the Kronecker product ($\otimes$) between the frequency-domain correlation ${\bf{R}}_{f}$ and the temporal-domain correlation ${\bf{R}}_{t}$ (21) implies that the frequency- and temporal-domain correlation structure is separable. This allows for a compact and computationally efficient representation of the channel covariance matrix. Building on this insight, we propose the following two-stage encoder design.
  • Frequency Encoder: In the Frequency Encoder stage, the pilot observations are processed along the subcarrier domain, extracting features on the frequency-domain correlation structure at pilot indices, corresponding to the rows of the matrix $\mathbf{R}_{f}\left[\mathcal{P},:\right]$.
  • Temporal Encoder: In the Temporal Encoder stage, the output of the Frequency Encoder is subsequently processed along the OFDM symbol domain, capturing temporal-domain correlation structures at pilot indices, corresponding to the rows of the matrix $\mathbf{R}_{t}\left[\mathcal{P},:\right]$.
This two-stage encoder enables the model to learn the joint frequency–temporal correlation structure; each stage uses a single encoder block, which is critical for the accurate MMSE estimation. Crucially, we note that this encoder design represents a key distinction of the proposed A-MMSE learning mechanism compared to conventional DNN-based methods. Specifically, existing approaches typically adopt standard learning architectures such as CNNs [36] or Attention Transformers [25], without explicitly tailoring them to the statistical properties inherent in OFDM channels. In contrast, our method explicitly incorporates the inherent temporal and frequency-domain correlation structures of OFDM channels into the learning process. This integration not only enhances learning efficiency but also improves the model’s capability to generalize under practical channel conditions. In this sense, the proposed A-MMSE approach is distinguished not only by its design, which enables the learning of a linear estimation filter rather than directly estimating the OFDM channel matrix (as illustrated in Fig. 1), but also by its advanced learning mechanism that explicitly reflects the inherent correlation structure of OFDM channels.
@@ IV-B2 Multi-Head Self-Attention Mechanism
Subsequently, the embedding matrix $\mathbf{X}_{\text{Emb}}$ is fed into the MHA module $\left(\mathcal{F}_{\text{MHA}}(\cdot)\right)$ with the number of heads fixed at $L$, where $L$ corresponds to the total number of DM-RS pilot symbols (one head per pilot). Each complete pilot is split into real and imaginary parts, resulting in the $2L$ column dimension of $\mathbf{X}_{\mathrm{Emb}}$. For example, with Type-A DM-RS and additional Position =1=1 (i.e., two DM-RS OFDM symbols) using a comb-2 pattern over 6 resource blocks (RBs), which has 72 subcarriers, there are $72/2=36$ pilots per DM-RS symbol, hence $L=36\times 2=72$ pilots in total. In general,
@@
where $N_{\mathrm{RB}}$ is the number of resource blocks, $N_{\mathrm{SC/RB}}=12$ is the number of subcarriers per RB, and $N_{\mathrm{DM-RS\,sym}}$ denotes the number of OFDM symbols carrying the DM-RS. The operation of the MHA module $\mathcal{F}_{\text{MHA}}(\cdot)$ is described in (12)–(15). We note that this allows each head to learn distinct frequency-domain correlation patterns. Following the MHA module, Add & Layer Normalization is applied, followed by a FFN, and subsequently another Add & Layer Normalization, in sequential order. Details on Add & Norm and FFN are provided in [37]. After these processes, the encoder extracts frequency-domain statistical information, which can be formally represented as
@@
where $\mathbf{Y}_{\text{Freq}}\in\mathbb{R}^{2L\times N}$ represents features encoded by the Frequency Encoder, $\mathcal{E}_{\text{Frequency}}$ represents the Frequency Encoder module, and $\theta_{\text{MHA, Freq}}$ and $\theta_{\text{Freq}}$ denote the learnable parameters of MHA module and the combined parameters of the two Add & Layer Normalization and FFN, respectively. Correspondingly, the overall operation $\mathbf{x}_{\text{in}}\mapsto\mathbf{Y}_{\text{Freq}}$ yields a real-valued matrix, where each row corresponds to a pilot position and columns align with subcarrier-specific embedding dimensions, explicitly capturing both local and global dependencies in the frequency-domain.
Temporal Encoder: In the second stage, a linear projection layer transforms the features encoded in the previous stage into a high-dimensional frequency and temporal domain, where the projection dimension is $NM$. The encoded feature matrix is given as $\mathbf{Y}_{\text{Freq}}=\left[\mathbf{y}_{\text{freq},1};\mathbf{y}_{\text{freq},2};\cdots;\mathbf{y}_{\text{freq},2L}\right]\in\mathbb{R}^{2L\times N}$. Here, each row vector $\mathbf{y}_{\text{Freq},i}$ is linearly projected as $\mathbf{x}_{\text{proj},i}=\mathbf{y}_{\text{Freq},i}\cdot\mathbf{P}_{\text{proj}}+\mathbf{b}_{\text{proj}}$, where $\mathbf{P}_{\text{proj}}\in\mathbb{R}^{N\times d_{p}}$ is a learnable projection matrix and $\mathbf{b}_{\text{proj}}\in\mathbb{R}^{d_{p}}$ is a learnable bias vector. Stacking all projected vectors $\mathbf{x}_{\text{proj},i}$, this yields the matrix given by
@@
Next, the projected matrix $\mathbf{X}_{\text{Proj}}$ is fed into another MHA module, with the number of heads fixed at $M$ (one head for each OFDM symbol). Similar to the above, this enables each head to effectively learn distinct temporal correlation patterns. The subsequent process is identical to that of the Frequency Encoder.
@@
where $\mathbf{Y}_{\text{Temp}}\in\mathbb{R}^{2L\times NM}$ represents features encoded by the Temporal Encoder, $\mathcal{E}_{\text{Temporal}}$ represents the Temporal Encoder module, with $\theta_{\text{MHA, Temp}}$ and $\theta_{\text{Temp}}$ being the learnable parameters of MHA module and the combined parameters of the two Add & Layer Normalization and FFN, respectively. The overall operations $\mathbf{Y}_{\text{Freq}}\mapsto\mathbf{Y}_{\text{Temp}}$ yield a real-valued matrix, whose rows correspond to pilot positions and whose columns align with subcarrier-symbol specific embedding dimensions. This explicitly captures both local and global dependencies in the frequency and temporal domain of the OFDM channel. The frequency-temporal features encoded by the two-stage encoder serve as inputs to the decoder.
@@
Once $\mathbf{Y}_{\text{Dec}}$ is obtained, it is partitioned into the real and imaginary components as follows.
@@ IV-B3 Training Method
where $N_{\mathcal{B}}$ denotes the number of mini-batches and $\mathcal{B}$ denotes the mini-batch size. In addition, $\mathbf{H}_{i,j}$ and $\hat{\bf{H}}_{i,j}$ represent the corresponding true channel matrix and estimated channel matrix, respectively. For the loss function $\mathcal{L}(\cdot)$, MSE or Huber loss can be chosen. We discuss this further in the next subsection. The loss function (31) is minimized via backpropagation and the Adam optimizer [17].